OpenStudy (anonymous):

If c = 30 and d = 8, then 2c + 4d =

OpenStudy (anonymous):

2*30+4*8=512

OpenStudy (anonymous):

Thanks but thats not a choice. A) 512 B) 278 C) 136 D) 92

OpenStudy (anonymous):

A. 512

OpenStudy (anonymous):

oh 512

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yes @Ella1

OpenStudy (anonymous):

not a problem:)

OpenStudy (anonymous):

i like your name

OpenStudy (anonymous):

Thank youu!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

How did you get 512 as an answer?

OpenStudy (anonymous):

2*30+4*8=512

OpenStudy (anonymous):

plug in the numbers 2(30)+4(8)= 512

OpenStudy (anonymous):

Human Calculator. lol

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

bye thanks for the help how old are you

OpenStudy (anonymous):

Your welcome. and im 17.

OpenStudy (anonymous):

am 16 my b-day is tomorrow

OpenStudy (anonymous):

Happy early Birthday!! :)

OpenStudy (anonymous):

Um isn't 2 + 30 = 60 and 4(8) = 32?

OpenStudy (anonymous):

Happy b-day @Ella1!

OpenStudy (anonymous):

thanks :) :)

OpenStudy (anonymous):

You have to multiply 2 and 30

OpenStudy (anonymous):